please factor completely thefollowing:
1.(x^2-y^2)
2.(x^2-2xy+y^2)
3. x^2+2xy+y^2

Answer by arunpaul(104)

1.(x^2-y^2) = (x+y)(x-y
2.(x^2-2xy+y^2)= (x-y)(x-y)
3. (x^2+2xy+y^2) = (x-y)(x-y)